OpenStudy (anonymous):

Help please!! log3(u^5⋅v^8). I am looking for another way to write this(simplified form). I was thinking maybe log3(u^5)(v^8) or log3(uv40) or maybe log3(u^5)(v(8). I really have no clue, any help appreciated.

OpenStudy (mertsj):

\[\log _{3}(u ^{5}v ^{8})=\log _{3}u ^{5}+\log _{3}v ^{8}=5\log _{3}u +8\log _{3}v\]

OpenStudy (anonymous):

Thank you